### v1

# From Language to Cognition: How LLMs Outgrow the Human Language Network

Abstract
Large language models (LLMs) exhibit remarkable similarity to neural activity in the human language network. However, the key properties of language shaping brain-like representations, and their evolution during training as a function of different tasks remain unclear. We here benchmark 34 training checkpoints spanning 300B tokens across 8 different model sizes to analyze how brain alignment relates to linguistic competence. Specifically, we find that brain alignment tracks the development of formal linguistic competence—i.e., knowledge of linguistic rules—more closely than functional linguistic competence. While functional competence, which involves world knowledge and reasoning, continues to develop throughout training, its relationship with brain alignment is weaker, suggesting that the human language network primarily encodes formal linguistic structure rather than broader cognitive functions. We further show that model size is not a reliable predictor of brain alignment when controlling for feature size and find that the correlation between next-word prediction, behavioral alignment and brain alignment fades once models surpass human language proficiency. Finally, using the largest set of rigorous neural language benchmarks to date, we show that language brain alignment benchmarks remain unsaturated, highlighting opportunities for improving future models. Taken together, our findings suggest that the human language network is best modeled by formal, rather than functional, aspects of language.
Keywords: Language; Human Language Network; LLMs; Brain Alignment; Behavioral Alignment
[…]
Introduction
Figure 1: Model Alignment with the Human Language Network is Primarily Driven by Formal Over Functional Linguistic Competence. (a) Average brain alignment across five Pythia models and five brain recording datasets, normalized by cross-subject consistency, throughout training. (b) Average normalized accuracy of the same models on formal linguistic competence benchmarks (two benchmarks). (c) Average normalized accuracy on functional linguistic competence benchmarks (six benchmarks). The results show that models first excel at formal linguistic tasks (such as judging the grammaticality of a sentence), aligning with the human language network, before later improving on functional benchmarks (such as reasoning about the world) but without further improvements in brain alignment. The x-axis is logarithmically spaced up to 16B tokens, capturing early training dynamics, and then evenly spaced every 20B tokens from 20B to  300B tokens. The vertical black line is at 16B tokens.
Deciphering the brain’s algorithms underlying our ability to process language and communicate is a core goal in neuroscience. Human language processing is supported by the brain’s language network (LN), a set of left-lateralized fronto-temporal regions in the brain (Binder et al., 1997; Bates et al., 2003; Gorno‐Tempini et al., 2004; Price, 2010; Fedorenko, 2014; Hagoort, 2019) that respond robustly and selectively to linguistic input (Fedorenko et al., 2024). Driven by recent advances in machine learning, large language models (LLMs) trained via next-word prediction on large corpora of text are now a particularly promising model family to capture the internal processes of the LN. In particular, when these models are exposed to the same linguistic stimuli (e.g., sentences or narratives) as human participants during neuroimaging and electrophysiology experiments, they account for a substantial portion of neural response variance (Schrimpf et al., 2021; Caucheteux and King, 2022; Goldstein et al., 2022; Pasquiou et al., 2022; Aw et al., 2023; Tuckute et al., 2024a; AlKhamissi et al., 2024; Rathi et al., 2025).
Dataset Modality Presentation Context Stimulus Example
Pereira2018 fMRI Reading Short Accordions produce sound with bellows …
Blank2014 fMRI Listening Long A clear and joyous day it was and out on the wide …
Fedorenko2016 ECoG Reading Short ‘ALEX’, ‘WAS’, ‘TIRED’, ‘SO’, ‘HE’, ‘TOOK’, …
Tuckute2024 fMRI Reading Short The judge spoke, breaking the silence.
Narratives fMRI Listening Long Okay so getting back to our story about uh Lucy …
Futrell2018 Reading Times Reading Long A clear and joyous day it was and out on the wide …
Table 1: Datasets Used for Evaluating Model Alignment. Neuroimaging datasets were collected via either functional magnetic resonance imaging (fMRI) or electrocorticography (ECoG). Text stimuli range from short sentences (Fedorenko2016, Tuckute2024) to paragraphs (Pereira2018) and entire stories (Blank2014, Narratives, Futrell2018) and were presented either visually or auditorily. Futrell2018 is a behavioral dataset.
Key Questions and Contributions
This work investigates four key questions, all aimed at distilling why LLM aligns to brain responses. Specifically, we investigate the full model development cycle as a combination of model architecture (structural priors) and how linguistic competence emerges across training (developmental experience). We ask: (1) What drives brain alignment in untrained models? (2) Is brain alignment primarily linked to formal or functional linguistic competence (Mahowald et al., 2024)? (3) Does model size or next-word prediction explain brain alignment? (4) Do current LLMs fully account for the explained variance in brain alignment benchmarks? To answer these questions, we introduce a rigorous brain-scoring framework to conduct a controlled and large-scale analysis of LLM brain alignment. Our findings reveal that the initial brain alignment of models with untrained parameters is driven by context integration in sequence-based models. During training, alignment primarily correlates with formal linguistic competence, which saturates early on in training ($\sim 4$B tokens), consistent with a plateauing of model-to-brain alignment. Functional competence, on the other hand, emerges later in training, tracks brain alignment less strongly, and continues to grow even when brain alignment has already saturated. This disconnect later in training is further exemplified by a fading of the correlation between models’ brain alignment and their next-word-prediction performance, as well as their behavioral alignment. Further, model size is not a reliable predictor of brain alignment when controlling for the number of features, challenging the assumption that larger models necessarily resemble the brain more. Finally, we show that current brain alignment benchmarks remain unsaturated, indicating that LLMs can still be improved in their ability to model human language processing.
Preliminaries & Related Work
A Primer on Language in the Human Brain
The human language network (LN) is a set of left-lateralized frontal and temporal brain regions supporting language. These regions are functionally defined by contrasting responses to language inputs over perceptually matched controls (e.g., lists of non-words) (Fedorenko et al., 2010). The language network exhibits remarkable selectivity for language processing compared to various non-linguistic inputs and tasks, such as music perception (Fedorenko et al., 2012; Chen et al., 2023) or arithmetic computation (Fedorenko et al., 2011; Monti et al., 2012) (for review, see (Fedorenko et al., 2024)) and the language network only shows weak responses when participants comprehend or articulate meaningless non-words (Fedorenko et al., 2010; Hu et al., 2023). This selectivity profile is supported by extensive neuroimaging research and further corroborated by behavioral evidence from aphasia studies: when brain damage is confined to language areas, individuals lose their linguistic abilities while retaining other skills, such as mathematics (Benn et al., 2013; Varley et al., 2005), general reasoning (Varley and Siegal, 2000), and theory of mind (Siegal and Varley, 2006).
Model-to-Brain Alignment
Prior work has shown that the internal representations of certain artificial neural networks resemble those in the brain. This alignment was initially observed in the domain of vision (Yamins et al., 2014; Khaligh-Razavi and Kriegeskorte, 2014; Cichy et al., 2016; Schrimpf et al., 2018, 2020; Cadena et al., 2019; Kubilius et al., 2019; Nayebi et al., 2018; Kietzmann et al., 2019; Zhuang et al., 2021) and has more recently been extended to auditory processing (Kell et al., 2018; Tuckute et al., 2023; Koumura et al., 2023) and language processing (Schrimpf et al., 2021; Caucheteux and King, 2022; Goldstein et al., 2022; Kauf et al., 2023; Hosseini et al., 2024; Aw et al., 2023; AlKhamissi et al., 2024; Tuckute et al., 2024b; Rathi et al., 2025).
Untrained Models
Recent work in vision neuroscience has shown that untrained convolutional networks can yield high brain alignment to recordings in the visual ventral stream without the need for experience-dependent training (Geiger et al., 2022; Kazemian et al., 2024). Other works have investigated the inductive biases in different architectures and initializations in models of visual processing (Cichy et al., 2016; Cadena et al., 2019; Geiger et al., 2022), speech perception (Millet and King, 2021; Tuckute et al., 2023), and language (Schrimpf et al., 2021; Pasquiou et al., 2022; Hosseini et al., 2024), highlighting that randomly initialized networks are not random functions (Teney et al., 2024).
Methods
Benchmarks for Brain Alignment
Neuroimaging & Behavioral Datasets
The neuroimaging datasets used in this work can be categorized along three dimensions: the imaging modality, the context length, and the modality through which the language stimulus was presented to human participants (auditory or visual). Table 1 provides an overview of all datasets in this study. We only consider neural units (electrodes, voxels, or regions) associated with the brain’s language network that were localized by their respective authors using the method described in the Functional Localization section, as packaged in Brain-Score Schrimpf et al. (2020, 2021). However, for the Narratives dataset, functional localization was not performed—to approximate the language regions, we leverage a probabilistic atlas for the human language system (Lipkin et al., 2022). Following the functional localization approach, we extract the top-10% language-selective voxels (from the probabilistic atlas) within anatomically defined language parcels. In an additional analysis, we investigate alignment with language behavior using the Futrell et al. (2018) dataset, which consists of self-paced per-word human reading times. See Appendix A for details of each dataset. To the best of our knowledge, this study examines the largest number of benchmarks compared to previous work, providing a more comprehensive and reliable foundation for understanding what drives brain alignment in LLMs. By analyzing diverse datasets, we ensure that our conclusions about brain-aligned models are not limited to specific stimuli or experimental conditions, leading to more generalizable insights.
Brain-Alignment Metrics
Following standard practice in measuring brain alignment, we train a ridge regression model to predict brain activity from model representations, using the same linguistic stimuli presented to human participants in neuroimaging studies (Schrimpf et al., 2020, 2021). We then measure the Pearson correlation between the predicted brain activations and the actual brain activations of human participants on a held-out set. This process is repeated over $k$ cross-validation splits, and we report the average (mean) Pearson correlation as our final result. We refer to this metric as Linear Predictivity. In the following section, we demonstrate why other metrics such as Centered Kernel Alignment (CKA; Kornblith et al., 2019) and Representational Similarity Analysis (RSA; Kriegeskorte et al., 2008) are not suitable measures for brain alignment on current language datasets.
Figure 2: Evaluating Brain Alignment with Linear Predictivity and No Contextualization is Most Stringent. (a) Average brain alignment across 8 Pythia models under three conditions: (1) a pretrained model processing the original stimuli, (2) a pretrained model processing random sequences of the same length (averaged over five random seeds) as a control condition, and (3) the model with untrained parameters processing the original stimuli. The linear predictivity metric differentiates between meaningful and random stimuli most strongly, while RSA and CKA overestimate alignment. (b) Brain alignment on the Pereira2018 dataset under two cross-validation schemes: with contextualization (random sentence split) and without contextualization (story-based split). Allowing contextualization makes it easier for untrained models to predict brain data, while enforcing stronger generalization via a story-based split reveals a significant drop in untrained model alignment.
Estimation of Cross-Subject Consistency
To assess the reliability of our datasets and account for the inherent noise in brain recordings, we compute a cross-subject consistency score (Feather et al., 2025), also referred to as the noise ceiling (Schrimpf et al., 2021). The consistency score is estimated by predicting the brain activity of a held-out subject using data from all other subjects, through 10-fold cross-validation of all subjects. To obtain a conservative ceiling estimate, we extrapolate subject pool sizes and report the final value based on extrapolation to infinitely many subjects. For Tuckute2024 we use the theoretical estimate provided by (Tuckute et al., 2024b). The consistency scores are provided in the Appendix. To aggregate scores across benchmarks, we normalize each model’s Pearson correlation ($r$) score for Linear Predictivity by the cross-subject consistency estimate, using the formula: (normalized score=raw $\textnormal{normalized score}=\frac{\textnormal{raw score}}{\textnormal{consistency}}$). The final brain alignment score for each model is reported as the average across all benchmarks. Otherwise, when reporting raw brain alignment, we compute the mean Pearson correlation across datasets without normalization.
Benchmarks for Formal & Functional Competence
There is substantial evidence in neuroscience research that formal and functional linguistic competence are governed by distinct neural mechanisms Mahowald et al. (2024). Formal linguistic competence pertains to the knowledge of linguistic rules and patterns, while functional linguistic competence involves using language to interpret and interact with the world. Therefore, to accurately track the evolution of each type of competence during training, we here focus on benchmarks that specifically target these cognitive mechanisms in LLMs.
Formal Linguistic Competence
To assess formal linguistic competence, we use two benchmarks: BLiMP (Warstadt et al., 2019) and SyntaxGym (Gauthier et al., 2020). BLiMP evaluates key grammatical phenomena in English through 67 tasks, each containing 1,000 minimal pairs designed to test specific contrasts in syntax, morphology, and semantics. Complementing this, SyntaxGym consists of 31 tasks that systematically measure the syntactic knowledge of language models. Together, these benchmarks provide a robust framework for evaluating how well LLMs acquire and apply linguistic rules.
Functional Linguistic Competence
Functional competence extends beyond linguistic rules, engaging a broader set of cognitive mechanisms. To assess this, we use six benchmarks covering world knowledge (ARC-Easy, ARC-Challenge (Clark et al., 2018)), social reasoning (Social IQa (Sap et al., 2019)), physical reasoning (PIQA (Bisk et al., 2019)), and commonsense reasoning (WinoGrande (Sakaguchi et al., 2019), HellaSwag (Zellers et al., 2019)). Together, these benchmarks provide a comprehensive evaluation of an LLM’s ability to reason, infer implicit knowledge, and navigate real-world contexts.
Metrics
We evaluate all benchmarks in a zero-shot setting, using surprisal as the evaluation metric—where the model’s prediction is determined by selecting the most probable candidate, as packaged in language model evaluation harness (Gao et al., 2024). We report accuracy normalized by chance performance, where 0% indicates performance at the random chance level.
Benchmark for Language Modeling
We use a subset of FineWebEdu Penedo et al. (2024) to evaluate the perplexity of the models on a held-out set. Specifically, use a maximum sequence length of 2048, and evaluate on the first 1000 documents of the AyCC-MAIN-2024-10 subset.
Large Language Models (LLMs)
Throughout this work, we use eight models from the Pythia model suite (Biderman et al., 2023), spanning a range of sizes: {14M, 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B}. Each model is evaluated across 34 training checkpoints, spanning approximately 300B tokens. These checkpoints include the untrained model, the final trained model, and 16 intermediate checkpoints that are logarithmically spaced up to 128B tokens. The remaining 14 checkpoints are evenly spaced every 20B tokens from 20B to 280B tokens, ensuring a comprehensive analysis of alignment trends throughout training. Since smaller models fail to surpass chance performance on many functional benchmarks, we exclude 14M, 70M, 160M from analyses that compare brain alignment with functional benchmark performance.
Rigorous brain-scoring
While substantial progress has been made in measuring alignment between LLM representations and neural activity, there remains no standardized approach for comparing results across different datasets and experimental conditions. Here, we aim to establish a set of desiderata for evaluating brain alignment. For a model to be considered truly brain-aligned, two key criteria must be met. First, high alignment scores should indicate that the model captures stimulus-driven responses—meaning that when presented with a random sequence of tokens, alignment should drop significantly compared to original linguistic stimuli. Second, a brain-aligned model should generalize effectively to new linguistic contexts rather than overfitting to specific examples. Below, we address these two points to justify our choice of metric and cross-validation scheme for each dataset. For all benchmarks, we localize language-selective units, which is consistent with neural site selection in neuroscience experiments and allows for fair comparisons across models irrespective of model size.
Figure 3: Context Integration drives Brain Alignment of Untrained Models. (a) Brain alignment of models with untrained parameters across five random initializations, showing that sequence-based models (GRU, LSTM, Transformers, and mean pooling) achieve higher alignment than models that rely solely on the last token representation (Linear, MLP), highlighting the importance of temporal integration. (b) Ablation study of architectural components in a single untrained Transformer-v2 block, demonstrating that attention mechanisms combined with positional encoding yield the highest brain alignment. (c) Diagram of the Transformer block architecture used in (b), with components grouped into attention (lower box) and MLP (upper box). (d) The average performance of five Pythia models with untrained parameters on formal and functional linguistic competence benchmarks, showing that formal competence exceeds chance level even in untrained parameter models.
Functional Localization
The human language network (LN) is defined functionally which means that units are chosen according to a ‘localizer’ experiment (Saxe et al., 2006). Specifically, the LN is the set of neural units (e.g., voxels/electrodes) that are more selective to sentences over a perceptually-matched control condition (Fedorenko et al., 2010). When selecting units from artificial models for comparison against LN units, previous work selected output units from an entire Transformer block based on brain alignment scores (Schrimpf et al., 2021). However, LLMs learn diverse concepts and behaviors during their considerable pretraining, not all of which are necessarily related to language processing, e.g., storage of knowledge (AlKhamissi et al., 2022) and the ability to perform complex reasoning (Huang and Chang, 2023). Therefore, we here follow the method proposed by AlKhamissi et al. (2024) that identifies language units in LLMs using functional localization as is already standard in neuroscience. This approach offers a key advantage: it enables direct comparisons across models by selecting a fixed set of units, identified through the independent localizer experiment. In this work, we localize $128$ units for all models unless otherwise specified, and we show in the Appendix that the results hold when selecting a different number of units.
Robust Metrics and Generalization Tests
Measuring Stimulus-Driven Responses
We first ask if the alignment procedure is meaningful, i.e., whether the encoding models capture meaningful linguistic information and generalize to new linguistic contexts. Figure 2(a) shows average brain alignment across all brain datasets under three conditions: (1) a pretrained model processing original stimuli, (2) a pretrained model processing random token sequences, and (3) an untrained model processing original stimuli. To evaluate metric reliability, we expect random sequences to yield significantly lower alignment than real stimuli. However, CKA fails this criterion, assigning similar alignment scores to both, and even untrained models surpass pretrained ones. In contrast, linear predictivity correctly differentiates between real and random stimuli, with a stronger separation than RSA.
Generalization and Contextualization
The second criterion we propose is that LLMs with high brain alignment should be able to generalize to held-out stimuli, with a preference for generalizing far outside the stimuli used for mapping model to brain activity. A key factor in designing a corresponding cross-validation scheme is contextualization—how the data is split into train and test sets Feghhi et al. (2024). The Pereira2018 dataset consists of 24 topics composed of multi-sentence passages, and sentences are presented in their original order to both humans and models. A random sentence split (contextualization) allows sentences from the same topic in both train and test sets, and is thus less demanding of generalization. A stronger generalization test ensures entire topics are held out, preventing models from leveraging shared context. Figure 2(b) shows that contextualization makes it easier for model to predict brain activity. In contrast, topic-based splits halve the raw alignment score for pre-trained models. The score of untrained models is reduced even more strongly when enforcing generalization across topics, suggesting that much of their alignment is context-dependent. Nonetheless, untrained models retain significant alignment – about half of pre-trained models – even with strong generalization requirements. We next explore the contributing factors for the alignment of models with untrained parameters.
Figure 4: Brain Alignment Saturates Early on in Training. Plots indicate the brain alignment scores of six models from the Pythia model suite with varying sizes (log x-axis up to 16B tokens, uneven spacing after black line). Scores are normalized by their cross-subject consistency scores. Alignment quickly peaks around 2–8B tokens before saturating or declining, regardless of model size.
What Makes Untrained Models Brain Aligned
In Figure 2 we show that untrained models, despite achieving lower alignment scores than their pretrained counterparts ($\sim 50\%$), still achieve relatively decent alignment and surpass that of the models evaluated with a random sequence of tokens. Therefore, we here ask, what are the main drivers for this surprising alignment.
Inductive Biases of Untrained Models
We evaluate the brain alignment of various LLMs with untrained parameters to determine which architecture exhibits the strongest inductive bias toward the human language network. Figure 3(a) presents the average alignment across five different random initializations for each untrained model. Each model consists of a stack of two building blocks from its respective architecture, with a hidden state of $1024$. To ensure a fair comparison, we apply the localizer to the output representations of the last token in the sequence from these two blocks, extracting 128 units to predict brain activity.
Our findings reveal two key insights. First, sequence-based models—such as GRU, LSTM, Transformers, and even a simple mean operation over token representations—exhibit higher brain alignment than models that rely solely on the last token’s representation, such as Linear or MLP. In other words, context or temporal integration is a crucial factor in achieving high alignment. Second, we observe a notable difference between Transformer-v1 and Transformer-v2. While Transformer-v2 applies static positional embeddings by directly adding them to token embeddings, Transformer-v1 uses rotary position encoding. Our results suggest that static positional encoding enables models to capture intrinsic temporal dynamics in sentences, providing further evidence that temporal integration is critical for brain-like language representations.
Figure 5: Formal Competence Tracks Brain Alignment More Closely Than Functional Competence. Each column compares how the evolution of formal competence (top) and functional competence (bottom) tracks the evolution of brain alignment during training. Formal and functional competence correspond to low-level and high-level linguistic processing, respectively. The $R^{2}$ values quantify the strength of this relationship, with higher values in formal competence suggesting it as the key driver of the observed brain alignment. (a): The data averaged across models of five different sizes. (b-d): the same comparison as in (a), but with comparisons were made for models from the Pythia suite with three different sizes.
Isolating Critical Components of Transformers
We perform an ablation study on the architectural components of Transformer-v2 using a single block to further isolate the key elements responsible for brain alignment of untrained parameter models (Figure 3(c)). To isolate the architectural components, we study the untrained model. The architectural components analyzed are labeled on the left of each bar in Figure 3(b). AyAttn refers to all components inside the lower box in Figure 3(c), including the first layer norm, multi-head attention, and the residual connection that follows. AyMLP corresponds to the components in the upper box, comprising the post-attention layer norm, MLP, and the subsequent residual layer. AyPos represents the addition of positional embeddings to token embeddings. AyTokens means the model directly returns the raw token embeddings without further processing. This structured ablation helps pinpoint the specific components that contribute most to brain alignment. Once again, we observe that integration across tokens, via attention mechanisms and positional encoding, yields the highest brain alignment. Further, we found that untrained parameter models perform better than chance-level performance on formal competence benchmarks, mirroring their non-zero brain alignment. In contrast, functional competence benchmarks remain at chance level for untrained models. This further supports the finding that brain alignment is primarily driven by formal, rather than functional, linguistic competence. (Figure 3(d)).
Role of Weight Initialization
Using the Transformer-v2 architecture from Figure 3(a), we varied the standard deviation used to initialize the weights of the untrained model (see Figure 9 in the Appendix). Our results reveal that the initialization value yielding the highest alignment is the default setting used in the HuggingFace Transformers library Wolf et al. (2019) suggesting that brain alignment may serve as a useful heuristic for optimizing model initialization. (see also Hosseini et al. (2024)).
How Does Alignment Change Over Training?
Having established what makes an untrained model brain-aligned in the previous section, we now investigate how brain alignment evolves during training. To do this, we use the Pythia model suite Biderman et al. (2023), which consists of models of various sizes, all trained on the same  300B tokens, with publicly available intermediate checkpoints for analysis.
Brain Alignment Evolution During Training
Figure 4 illustrates the brain alignment of six Pythia models across five brain recording datasets at 34 training checkpoints, spanning approximately 300B tokens. Each panel presents checkpoints that are logarithmically spaced up to the vertical line, emphasizing the early-stage increase in brain alignment, which occurs within the first 5.6% of training time. Beyond this point, the panels display the remaining training period, where brain alignment stabilizes. More specifically, we observe the following trend: (1) Brain alignment is similar to the untrained model until approximately 128M tokens. (2) A sharp increase follows, peaking around 8B tokens. (3) Brain alignment then saturates for the remainder of training. Despite the vast difference in model scales shown in Figure 4, the trajectory of brain alignment is remarkably similar.
Brain Alignment Tracks Formal Competence
Following the observation that brain alignment plateaus early in training, we next investigate how this relates to the development of formal and functional linguistic competence in LLMs. Figure 5 displays the average brain alignment alongside the average performance on formal competence benchmarks (top row) and functional competence benchmarks (bottom row). This is shown for three Pythia models (1B, 2.8B, and 6.9B parameters) and the average of five Pythia models (last column) across the training process. To quantify this relationship, we trained a ridge regression model (with a single scalar weight) to predict brain alignment scores from benchmark scores using 10-fold cross-validation. The average R-squared value across these folds serves as our metric for comparing the relationship between formal/functional linguistic competence and brain alignment. These R-squared values are shown in each panel of Figure 5. Finally, we performed a Wilcoxon signed-rank test on the distributions of R-squared values. This revealed that formal competence is significantly more strongly correlated with brain alignment than functional linguistic competence (W = 0.0, p $<$ 0.002). One possible explanation for why brain alignment emerges before formal linguistic competence is that existing LLM benchmarks assess performance using discrete accuracy thresholds (hard metrics), rather than capturing the gradual progression of competence through more nuanced, continuous measures (soft metrics) (Schaeffer et al., 2023).
We show the individual benchmark scores across all checkpoints in Figure 8.
Figure 6: NWP and behavioral alignment correlate with brain alignment only in early training. (Top Row): Correlation between brain alignment and language modeling loss shows a strong, significant relationship during early training (up to 2B tokens). While this correlation weakens in later stages (up to  300B tokens). Results are shown for three models and the average of all 8 models (last column). (Bottom Row): The same analysis, but for the correlation between brain alignment and behavioral alignment, revealing a similar trend—strong correlation early in training, but no significant relationship as models surpass human proficiency.
Superhuman LLMs Lose Behavioral Alignment
Human language processing is strongly modulated by prediction, with words that violate predictions leading to longer reading times (Smith and Levy, 2013). Early in training, LLMs align with this pattern, but as they surpass human proficiency (Shlegeris et al., 2022), their perplexity drops and they begin encoding statistical regularities that diverge from human intuition (Oh and Schuler, 2023; Steuer et al., 2023). This shift correlates with a decline in behavioral alignment, suggesting that superhuman models rely on different mechanisms than those underlying human language comprehension. Figure 6 shows that brain alignment initially correlates with perplexity and behavioral alignment, but only during the early stages of training (up to  2B tokens). Beyond this point, as models surpass human language proficiency, these correlations diminish. In larger models, we observe a negative correlation between brain alignment and behavioral alignment in the later stages of training. This trend reinforces the idea that early training aligns LLMs with human-like processing as also observed in earlier, while in later stages their language mechanisms diverge from humans.
Model Size Does Not Predict Alignment
Figure 7: Model Size Does Not Predict Brain Alignment when localizing a fixed set of language units. Brain alignment across model sizes in the Pythia suite, measured at their final training checkpoints. Brain alignment is shown for each dataset, along with the average alignment across datasets, for eight models of varying sizes.
Figure 7 presents the brain alignment for each dataset, along with the average alignment across datasets, for eight models of varying sizes from the Pythia model suite (final checkpoint). Contrary to the assumption that larger models exhibit higher brain alignment Aw et al. (2023), we observe a decline in average alignment starting from 1B parameters up to 6.9B parameters, when controlling for feature size. This analysis is made possible by functional localization, which allows us to extract a fixed number of units from each model, rather than relying on hidden state dimensions, as done in previous studies. This approach ensures a fairer comparison among models. We show in the Appendix that increasing the number of localized units has minimal impact on the relative ranking of the models. Additionally, these findings align with expectations in the neuroscience language community, where it is widely believed that human language processing does not require superhuman-scale language models to capture neural activity in the brain’s language network.
Conclusion & Future Work
In this work, we investigated how brain alignment in LLMs evolves throughout training, revealing different learning processes at play. We demonstrated that alignment with the human language network (LN) primarily correlates with formal linguistic competence Mahowald et al. (2024), peaking and saturating early in training. In contrast, functional linguistic competence, which involves world knowledge and reasoning, continues to grow beyond this stage. These findings suggest that the LN primarily encodes syntactic and compositional structure, in line with the literature of language neuroscience Fedorenko et al. (2024), while broader linguistic functions may rely on other cognitive systems beyond the LN. We also showed that model size is not a reliable predictor of brain alignment when controlling for feature size. Instead, alignment is determined by architectural inductive biases, token integration mechanisms, and training dynamics. Our standardized brain-scoring framework eliminates contextualization biases from previous work, ensuring more rigorous evaluations. Finally, we demonstrated that current brain alignment benchmarks are not saturated, indicating that LLMs can still be improved in modeling human language processing. Together, these findings challenge prior assumptions about how alignment emerges in LLMs and provide new insights into the relationship between artificial and biological language processing.
Future Work
While this study offers a comprehensive analysis of brain alignment in LLMs, several open questions remain. If functional competence extends beyond the language network, future work should explore which additional brain regions LLMs align with as they develop reasoning and world knowledge, particularly in other cognitive networks like the multiple demand (Duncan and Owen, 2000) or theory of mind network (Saxe and Kanwisher, 2003; Saxe and Powell, 2006). Our findings suggest that LLM brain alignment studied should be broadened from the LN to downstream representations underlying other parts of cognition. This raises the question of whether specific transformer units specialize in formal vs. functional linguistic competence (AlKhamissi et al., 2024; Kumar et al., 2024). A key question remains: Does LLM alignment evolution mirror human language acquisition? Comparing it to developmental data could reveal insights into their learning trajectories, and help differentiate formal from functional language learning. Expanding brain-scoring benchmarks and incorporating multimodal models will help address these questions, further bridging the gap between artificial and biological intelligence and deepening our understanding of how both systems process and represent language.
Appendix
Appendix A Neuroimaging & Behavioral Datasets
Neuroimaging Datasets
Pereira et al. (2018)
This dataset consists of fMRI activations (blood-oxygen-level-dependent; BOLD responses) recorded as participants read short passages presented one sentence at a time for 4 s. The dataset is composed of two distinct experiments: one with 9 subjects presented with 384 sentences, and another with 6 subjects presented with 243 sentences each. The passages in each experiment spanned 24 different topics. The results reported for this dataset are the average alignment across both experiments after normalizing with their respective cross-subject consistency estimates.
Blank et al. (2014)
This dataset also involves fMRI signals but recorded from only 12 functional regions of interest (fROI) instead of the higher resolution signal used by Pereira et al. (2018). The data was collected from 5 participants as they listened to 8 long naturalistic stories that were adapted from existing fairy tales and short stories (Futrell et al., 2018). Each story was approximately 5 minutes long, averaging up to 165 sentences, providing a much longer context length than the other neuroimaging datasets. When measuring brain alignment, we use the input stimuli of the last 32 TRs as the model’s context.
Fedorenko et al. (2016)
This dataset captures ECoG signals from 5 participants as they read 8-word-long sentences presented one word at a time for 450 or 700 ms. Following Schrimpf et al. (2021) we select the 52/80 sentences that were presented to all participants.
Tuckute et al. (2024b)
In this dataset, 5 participants read 1000 6-word sentences presented one sentence at a time for 2 s. BOLD responses from voxels in the language network were averaged within each participant and then across participants to yield an overall average language network response to each sentence. The stimuli used span a large part of the linguistic space, enabling model-brain comparisons across a wide range of single sentences. Sentence presentation order was randomized across participants. In combination with the diversity in linguistic materials, this dataset presents a particularly challenging dataset for model evaluation.
Narratives Dataset (Nastase et al., 2021)
This dataset consists of fMRI data collected while human subjects listened to 27 diverse spoken story stimuli. The collection includes 345 subjects, 891 functional scans, and approximately 4.6 hours of unique audio stimuli. For our story-based analysis, we focused on 5 participants who each listened to both the Lucy and Tunnel stories. Since functional localization was not performed in the Narratives dataset, we approximated language regions by extracting the top-10% voxels from each anatomically defined language region according to a probabilistic atlas for the human language system (Lipkin et al., 2022). Due to the limited corpus of two stories, traditional 10-fold cross-validation was not feasible. To implement topic-based splitting while maintaining methodological rigor, we partitioned each story into $n$ distinct segments, with each segment functioning as an independent narrative unit. This segmentation approach effectively prevented cross-contamination of contextual information between splits, thereby preserving the integrity of our evaluation framework.
Appendix B Brain-Score Using Additional Metrics
Metrics: CKA & RSA
Centered Kernel Alignment (CKA)
Kornblith et al. (2019) introduced CKA as a substitute for Canonical Correlation Analysis (CCA) to assess the similarity between neural network representations. Unlike linear predictivity, it is a non-parameteric metric and therefore does not require any additional training. CKA is particularly effective with high-dimensional representations, and its reliability in identifying correspondences between representations in networks trained from different initializations (Kornblith et al., 2019).
Representational Similarity Analysis (RSA)
Kriegeskorte et al. (2008) introduced RDMs as a solution to the challenge of integrating brain-activity measurements, behavioral observations, and computational models in systems neuroscience. RDMs are part of a broader analytical framework referred to as representational similarity analysis (RSA). In practical terms, to compute the dissimilarity matrix for an $N$-dimensional network’s responses to $M$ different stimuli, an $M$×$M$ matrix of distances between all pairs of evoked responses is generated for both brain activity and the language model’s activations Harvey et al. (2023). The correlation between these two matrices is then used as a measure of brain alignment.
Appendix C Additional Analysis on Untrained Models
Figure 9: Role of weight initialization on brain alignment in untrained models The default initialization standard deviation in the HuggingFace library (sd = 0.02) yields the highest brain alignment for untrained models, suggesting that initialization choices play a crucial role in shaping alignment even before training begins.
Figure 9 examines the effect of weight initialization variance on brain alignment in untrained models. We systematically vary the initialization standard deviation (sd) and find that the default HuggingFace initialization (sd = 0.02) achieves the highest alignment across datasets. This suggests that even before training begins, the choice of initialization can significantly influence how well a model’s representations align with neural activity. This finding raises an intriguing hypothesis: could brain alignment, a computationally inexpensive metric, serve as a useful heuristic for selecting optimal initialization parameters? If so, it could help models learn tasks more efficiently and converge faster, reducing the need for extensive trial-and-error in training from scratch.
Appendix D Formal & Functional Benchmark Scores
Figure 8 presents the individual benchmark scores for both formal and functional linguistic competence across training. Formal benchmarks peak early, mirroring the trajectory of brain alignment, and remain saturated throughout training. In contrast, functional benchmarks continue to improve, reflecting the models’ increasing ability to acquire factual knowledge and reasoning skills as they are trained on significantly more tokens using next-word prediction.
Appendix E Effect of Number of Units on Brain Alignment
Figure 10: The effect of the number of localized units on final brain alignment Brain alignment is evaluated after localizing 128, 1024, and 4096 units. While increasing the number of units slightly affects overall alignment, the relative ranking of models remains largely unchanged, indicating that model comparisons are robust to the choice of unit count.
Figure 10 illustrates the impact of localizing more units on final brain alignment across the eight Pythia models used in this study. We find that increasing the number of units has minimal impact on the relative ranking of models, with only a slight increase in average alignment. Additionally, model size does not influence brain alignment once the number of units is controlled, reinforcing the idea that alignment is driven by feature selection rather than scale.
Appendix F Cross-Subject Consistency Scores
Benchmark Consistency Score
Pereira2018 (Exp 2) 0.086
Pereira2018 (Exp 3) 0.144
Blank2014 0.178
Fedorenko2016 0.222
Tucktue2024 0.559
Narratives 0.181
Futrell2018 0.858
Table 2: Cross-subject consistency scores The values used to normalize the raw Pearson correlation. Pereira2018 (Exp 2) was computed without extrapolation.
Table 2 shows the cross-subject consistency scores computed with extrapolation for the different benchmarks used in this work.